If I add 55 mL of water to 125 mL of a 0.15 M NaOH solution, what will the new molarity of the diluted solution be?
1
Expert's answer
2020-04-22T10:54:19-0400

The new molarity is:

C = n/V = 0.125*0.15/(0.125 + 0.055) = 0.104 M